OpenStudy (anonymous):

compute modulus and argument a.1+i b.4-4i c.2+5i

OpenStudy (alexwee123):

argument is arctan(y/x) for a complex number x + iy

OpenStudy (anonymous):

1+i = sqrt 2 (modulus) 4-4i = sqrt(4^2+4^2) = sqrt32

OpenStudy (alexwee123):